UPDATE: 14th April 20:45

Once again, please accept our apologies for the problems you’ve seen over the previous couple of days. We realise that this has caused you all serious issues and for that, we’re very sorry.

Various internal changes have been implemented over the last 48 hours and currently, we believe that the network and associated services are now stable and will remain that way. We continue to monitor the situation closely to ensure that our network remains stable and there’s no further impact to your services.

Please email us in the normal way if you have any question or concerns. Thanks again for your support through this incident.

Ongoing DDoS attack against our network.

23:30 Again, our apologies

We continue to undertake remedial work to mitigate this ongoing attack.

We will update here as usual.

————

We are currently seeing a new, large-scale DDoS attack against our IP range. We are working to mitigate this but some services are being affected, with packet loss, routing failures or intermittent outages. Some email delivery will be queued until this is resolved.

We will update here as usual.

13:06 UPDATE

We are mitigating a large portion of this attack traffic but currently, the transit links remain saturated which is causing the current ongoing problems. We continue to work to resolve this as quickly as possible & apologise for the ongoing inconvenience caused.

13:44 UPDATE

We are seeing most services recovering. The attack target remains offline but we believe that this incident is now contained. We apologise again for this interruption in service. If you are still seeing issues, please restart your equipment. Tickets can be raised now in the normal manner & the support line remains ready to assist.

17:43 UPDATE

The offsite server that hosts the status NOC site went down during the afternoon. Purely coincidental, but it meant we weren’t able to access it to add more frequent updates. It’s now back and we’ll update the status on the DDoS attack issues shortly. Our apologies that this wasn’t available when it was needed the most.

10th April 2021 – internet issues [update]

We are currently seeing a large scale DDoS attack against our IP range.

This will lead to significant packet loss and access issues to our customers. Our NOC team are already at work to mitigate this. We will post a further update as soon as we have it

[Update 11/04/21 – 10:00am]
We believe the issue cleared shortly after 7pm yesterday. We are still monitoring this closely however we do not believe there is currently any ongoing customer impact

ISSUE: Virgin Media circuits – RESOLVED

We have seen alerts that some circuits from Virgin Media have dropped. We believe this is a Virgin Media issue but are currently investigating.
 
We will update this status as soon as we know more and within the hour at the latest.
 
UPDATE: 16:09
Virgin have confirmed an issue at one of their hub sites affecting some parts of Hertfordshire. Senior engineers and the Core Incident team are working on this at the moment.
 
We will have another update — if it’s not restored — within approx 30 minutes.
 
UPDATE: 16:20
We are seeing the affected circuits back on line – at this point we don’t have an ‘all clear’ so they should be seen at risk – but we hope this is resolved
 
RESOLVED 16:20
We have received the following update from Virgin Media
Please be advised that @ 16:22
Restoration Details:IOM Card 2 and MDA Card 1 remotely reseated on T0090 Luton Metnet 2a.
restoring all services.
 
This appears to have been a card failure at one of the Virgin Media Hub Sites – we apologise for the issues that occured here

Network Instability [resolved]

We are aware of some instability on our core London network – this has caused a number of short periods of packet loss this afternoon. Our NOC team are investigating this currently. We will update this as we know more.

This is being worked on as a high  priority currently

We apologise for any issues this may cause

[update 19:10]

Investigations are ongoing on the cause of this issue. Our engineers continue to work to understand and resolve the root cause of this issue. We can see BGP sessions drop between our core routers but do not fully understand the actual cause of this. Many users may not notice the issue however there will be occasional packet loss and ‘strange’ routes while this is being worked on

[Update 20:30]

While this issue has improved this is still ongoing. We are seeing BGP between two of our core routers (Telehouse North and East) flapping. This is causing some packet loss / routing drops. now we have isolated the reason for the drops  we are working to find why this is ongoing. Initial thoughts where that the routers had stale data and we therefore rebooted them – However the issue while improved has not gone away. We will update this further shortly

[Update 23:00]

The network has now been stable for over an hour and we believe the issue is cleared. we will however monitor this overnight before we close this. It appears that this was down to a forwarding issue on a core Juniper Switch in Telehouse West. After ruling out other (more likely) causes, this switch was reloaded approx 22:00 and so far the routing issues and flaps have calmed.

As the switches showed no obvious errors or log issues we will not close this until we are certain – and will investigate in more detail. However we hope there is no further customer issues as a result of this. We are sorry if this issue affected your service today

[Update 26/4/202 10:00]

The network has been stable overnight following the reboot of the core Juniper Switch. We are continuing to monitor – however we do believe this issue is now resolved.

We are sorry for any issues caused by this problem

UPDATE: Network outage 3rd Feb 2020 [resolved]

4th Feb: the supplier network resolved all oustabnding issues on this extended outage. Apologies again to anyone affected.

10.35AM UPDATE: this problem which affects a large number of FTTx lines is still being worked on by Openreach engineers. Latest estimate for the work to be completed is 18.30 tonight. We’ll continue to update here as and when we hear anymore news.

———

We are aware that one of our suppliers has a major service outage. The supplier engineers have no ETA to fix as yet. We will; update as we hear more. This will affect ADSL & FTTC lines. Our apologies for this outage.

 

Line affected:

Start: 03/02/2020 04:15 Next update / Cleared: 03/02/2020 09:30 (estimated)
Raised: 03/02/2020 04:15 Cleared Reason: N/A
Detected: 03/02/2020 04:15 Exchange Name: KINGS LANGLEY, LITTLE GADDESDEN, BOLSOVER, STAVELEY, HOLMEWOOD, OLD WHITTINGTON, CLAY CROSS, HOLYMOORSIDE, BASLOW, STONY STRATFORD, HANSLOPE, YARDLEY GOBION, CROXTON, CAXTON, LAISTERDYKE, DUDLEY HILL, UNDERCLIFFE, GRANTON, SOWERBY BRIDGE, ILLINGWORTH, CALDER VALLEY, WOMBWELL,
Incident Headline: Service affecting outage – Loss of Service
Incident Details: Our supplier engineers are working on the fault.
Area Codes: 01131 01132 01133 01138 01162 01173 01179 01212 01213 01214 01215 01216 01217 01223 01226 01246 01257 01274 01275 01311 01312 01313 01314 01315 01316 01317 01332 01422 01442 01446 01454 01480 01509 01773 01827 01908 01923 01926 01942 01954 02010 02032 02033 02081 02082 02083 02084 02085 02086 02087 02089 02476 02911 02920 02921 02922